OpenStudy (anonymous):

Convert the mixed number to an improper fraction. 3and5/6

OpenStudy (anonymous):

@Missiey

OpenStudy (alexandervonhumboldt2):

in main form: \[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\]

OpenStudy (alexandervonhumboldt2):

\[3\frac{ 5 }{ 6 }=\frac{ 3*6+5 }{ 6 }=simplify\]

OpenStudy (anonymous):

23/6

OpenStudy (anonymous):

@AlexandervonHumboldt2 is that the answer